OpenStudy (anonymous):

Solve the square root of the quantity of 14 + 2 plus 4 equals 8 and determine if true.

OpenStudy (anonymous):

\[\sqrt{14+2}+4=8\]??

OpenStudy (anonymous):

Yes that's it

OpenStudy (anonymous):

step by step what is \(14+2\)?

OpenStudy (anonymous):

16 lol

OpenStudy (anonymous):

and \(\sqrt{16}=?\)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

and \(4+4=?\)

OpenStudy (anonymous):

8

OpenStudy (anonymous):

so true i guess right?

OpenStudy (anonymous):

Alright I got it, thanks

OpenStudy (anonymous):

yw